Obsession Inc. is a publicly traded company with two business lines: perfumes and apparel. The beta of this company is 1.72. Obsession Inc. is contemplating a major restructuring in which it wants to sell off its apparel division, which has an unlevered beta of 0.71, for 20 million. It also wants to borrow an additional 25 million and buy back stock worth 45 million. After the sale of the division and the share repurchase, Obsession Inc. will have 46 million in debt and 143 million in equity outstanding. Assume a tax rate of 40.00% percent. The value of debt and equity prior to these changes are:

46 mn and 143 mn

21 mn and 188 mn

71 mn and 143 mn

46 mn and 188 mn

71 mn and 143 mn
Answer #1

After borrowing 25 million for repurchase ,value of debt=46 million

It means before borrowing ,value of debt=46-25=21 million

Now,After repurchase of stock of 45 million,Equity =143 million

Hence before repurchase,Equity=143+45=188 million

Hence, value of debt and equity prior to changes =

21 mn and 188 mn